OpenStudy (anonymous):

You are desining a patio . Your plans use a scale of 1/8 inch to 1 foot. The patio is 24 feet by 36 feet . What are its dimensions on the plans

hero (hero):

Hint: Use the following proportions: \(\large\frac{1 \text{ft}}{24 \text{ft}} = \frac{.125 \text{in}}{x \space \text{in}}\)

hero (hero):

\(\large\frac{1 \space \text{ft}}{36 \space \text{ft}} = \frac{.125 \space \text{in}}{x \space \text{in}}\)